# DoVer: Intervention-Driven Auto Debugging for LLM Multi-Agent Systems

### v2

@@ 1 Introduction
The advancement of Large Language Models (LLMs) has led to the rapid rise of LLM-based agent systems, particularly multi-agent architectures where agents of different roles work collaboratively to solve complex tasks Li2023camel; Gao2023autogen; MetaGPT. As these systems are increasingly developed and deployed in production, the need to debug their failures becomes inevitable during their lifecycle. Importantly, by “failures” we do not refer to conventional software errors (e.g., exceptions or crashes), but rather to the errors where the system executes without interruption yet produces incorrect or unsatisfactory results GAIA; AssistantBench. Such failures frequently arise in scenarios where one diagnoses why an agent system underperforms on benchmark tasks during the development phase, or when one addresses user-reported dissatisfaction (e.g., a ‘thumbs-down’ signal with textual feedback) from an online deployed system.
Debugging failures in LLM-based agent systems presents unique challenges. These systems typically involve multiple rounds of LLM calls, each with extensive textual context, making manual log inspection labor-intensive. Furthermore, in multi-agent tasks, tracking inter-agent information flow is crucial, as failures often stem from Inter-Agent Misalignment MAST. Recent efforts address these issues by using LLMs to analyze system failures DevAI; whoandwhen; AgentTracer, often via single-agent, single-step failure attribution. In this method, an execution log is input to an LLM tasked with identifying the agent and step responsible for the failure whoandwhen. However, as shown in our reproduction study (Section 3), log-based failure attribution is fundamentally limited by the uncertainty of ground-truth annotations. This uncertainty arises for several reasons: agent systems often employ multiple strategies (e.g., ReAct React) with distinct failure points in a single session, and inter-agent misalignment can render the assignment of responsibility to a single agent or step ambiguous.
@@
DoVer also supports interventions across multiple steps rather than restricting to single-point edits. By decomposing the failure trace into separate trials, we intervene at each and assess the impact. Within the Magnetic-One (M1) Magentic-One agent framework, our experiments show that DoVer recovers 18% and 28% of failures on datasets from AssistantBench AssistantBench and GAIA GAIA, respectively. It further enables the validation or refutation of 30–60% of failure hypotheses, depending on task complexity. To demonstrate the generality of DoVer, we apply it to another agent framework, AG2 autogen02, using a different dataset, GSMPlus li2024gsm. DoVer again performs effectively, achieving a 49% flip rate on failure cases.
@@ 2 Related Work
LLM-based agent systems exhibit diverse and frequent failure patterns that accumulate along long execution logs. To characterize why and where errors arise, MAST MAST catalogs failures across task interpretation, planning, tool/environment interaction, and verification, while evaluation frameworks DevAI; AgentEval argue that end-to-end pass/fail is too coarse and introduce requirement graphs or task-utility criteria that reveal where progress stalls. Extending this perspective to execution logs, TRAIL TRAIL creates turn-level traces and a fine-grained taxonomy (reasoning, planning, execution), empirically showing that even strong long-context models struggle at trace debugging.
A parallel line of work seeks failure attribution: identifying the earliest decisive step or agent that is responsible for the earliest sufficient cause of failure whoandwhen; AgentTracer. However, these attributions are inferred from logs and remain an untested hypothesis unless validated by execution. In DoVer, we treat attribution as a hypothesis to be tested. We apply a targeted edit at the implicated location (message, plan, tool call), and rerun the system, judging success by milestone/utility gains. This places emphasis on verified repair and is consistent with trajectory-aware evaluation advocated in AgentEval; TRAIL.
Recent contemporaneous studies [^1] These works appeared during the final stage of this work preparation or the subsequent peer-review period. further advance failure attribution by introducing reasoning-driven judges zhuRAFFLESReasoningbasedAttribution2025, abduct–act–predict counterfactual scaffolding westAbductActPredict2025, causal-inference formulations maAutomaticFailureAttribution2025, spectrum-based failure attribution approach geWhoIntroducingFailure2025, hierarchical error-attribution method banerjee2025didwronghierarchicallook, and graph-guided failure tracing approach zhang2025graphtracergraphguidedfailuretracing. Notably, the observation in westAbductActPredict2025 that adding explicit step indices improve attribution aligns with our own prompt-refinement study in Section 3.
Emerging works have also explored specialized failure-tracer models trained on curated success/failure trajectories AgentTracer; Aegis. These approaches are orthogonal to DoVer and could be incorporated into the initial failure-attribution stage of DoVer to strengthen its diagnostic signal.
@@ 2.2 Debugging Approaches for LLM-based Agent Systems
Beyond failure analysis and attribution, several systems explore how to debug trajectories, i.e., intervention and replay. Human-in-the-loop tools such as AGDebugger AGDebuger enable rewind/edit/re-execute with trace visualization, and graph runtimes like LangGraph langgraph-time-travel provide checkpoints, interrupts, and “time-travel” branching. These demonstrate that small, targeted interventions often work, but they are manual and hard to scale. The recent AgentDebug zhuWhereLLMAgents2025 work employs an intervention-driven methodology similar to DoVer, although it does not place particular emphasis on multi-agent settings.
From the software-repair perspective, AgentIssue-Bench package real agent-system issues into executable environments with failing tests (AgentIssue-Bench) and find low resolution rates for current software engineering agents, highlighting the difficulty of maintaining agent software. Industrial experience at Google similarly evaluates agent-based program repair on production bugs, showing promise but also current limits SE-Repair. This motivates auto debugging that verifies edits within the original run and measures intermediate progress. DoVer does so by choosing a minimal intervention, re-running the trajectory, and scoring milestone/utility gains, in line with DevAI; TRAIL.
@@ 3 From Log-Based Attribution to Intervention-Based Debugging
A common setup for log-based failure attribution proceeds as follows. The input is a session log, typically from an LLM-driven multi-agent system, covering the full trajectory from an initial user query to either a final answer or termination caused by system-defined stopping rules (e.g., a maximum number of replanning rounds). The log is a sequence of agent messages organized by turns. [^2] We do not consider logs from asynchronous agent executions in this work. The task is to identify the earliest single agent and single step responsible for the failure. A failure step is defined as a decisive error: if one were to replace the agent’s incorrect action at that step with a correct one, the trajectory would subsequently succeed. To evaluate this setup, the Who&When (WW) dataset was introduced in whoandwhen. In WW, multiple annotators independently reviewed session logs and then reconciled discrepancies through discussion. Reported results show that state-of-the-art LLMs at the time achieved below 10% accuracy on failure step attribution.
@@
Figure 1: Failure trace of Case 3 in WW-HC, illustrating ambiguity in failure attribution. The session consists of four distinct trials, each initiated by a plan update and executed via a ReAct-style React loop. Different strategies (e.g., direct scrolling in Trial 1 vs. calendar navigation in Trial 2) yield separate error points, making single-step attribution across the session inherently ambiguous. Trial 2 (Steps 53–55) further shows inter-agent misalignment: the Orchestrator issued an invalid instruction, while the WebSurfer compounded the error by executing an unrelated action.
@@
(1) Multiple trials within a single session. Modern agentic systems frequently employ ReAct-style React loops with repeated planning–execution cycles, producing multiple trials of task solving. We define a trial as the contiguous span starting from a planning step and continuing through the execution steps for that plan. Each trial may contain its own decisive error step(s), especially when exploring new strategies or branches. For instance, Figure 1 illustrates the failure trace of Case 3 in WW-HC, which comprises four distinct trials within a single session. Different strategies were explored (e.g., Trial 1 attempted to locate the target webpage via direct scrolling, whereas Trial 2 used the calendar feature), each introducing distinct potential error points. Consequently, enforcing a single-step attribution across the entire session is intrinsically ambiguous. In our review, 9 of the 14 uncertain cases exhibit this pattern.
(2) Ambiguity from multi-agent coordination. Attribution to a specific agent/step can be unclear when the underlying issue stems from inter-agent coordination. For instance, when a sub-agent fails to carry out an instruction, responsibility may lie with the sub-agent’s capabilities or with ambiguous/misaligned guidance from a supervising agent. In Trial 2 of Figure 1, for example, the Orchestrator agent instructed the WebSurfer agent to click on a non-existent control. Rather than reporting the issue, the WebSurfer instead clicked on an unrelated control that had no connection to the target year “2015”. In this case, both agents exhibited failures, making attribution to either one alone inappropriate. Such ambiguous attribution aligns with the Inter-Agent Misalignment category identified in MAST. We observed this phenomenon in 5 of the 14 uncertain cases.
@@ 4.1 DoVer Pipeline
Trial Segmentation. After a task is executed by an agentic system, we have obtained a long execution session log $\tau=\{(a_{t},m_{t},\sigma_{t})\}_{t=1}^{T}$, where $a_{t}$ represents the active agent that produces the message $m_{t}$ at step $t$, and $\sigma_{t}$ keeps stateful information that is necessary for state restore and replay. For example, this includes historical context (e.g., prior messages sent to LLMs) as well as browsing history for agents acting in a web-browsing role. Since modern LLM-based agent systems Magentic-One; smolagents; autogen02 have the self-reflection capabilities React; Reflexion, and re-plan after reflection, we then segment the trace $\tau$ into trials $\tau^{i}$} using re-plan steps as segmentation point, as illustrated in Figure 2. This trial segmentation shortens context so LLMs can reason about a single causal chain, and enables independent and parallel interventions for efficiency. Thus, multiple hypotheses can be proposed for a single session trace, which is essential as many failures admit more than one viable repair.
@@ 4.2 Evaluation Metrics
Specifically, when an intervention does not yield full success and human-annotated solution steps are available (e.g., in the GAIA benchmark GAIA), we could evaluate the degree of progress by comparing the new execution trace against human-annotated steps. For each log $\tau$, we extract up to $K\leq 5$ milestones $\{\mathbf{m}^{k}\}_{k=1}^{K}$ and measure how many of these milestones the new trace accomplishes. Both the extraction of milestones and their evaluation are performed using LLMs with carefully designed instructions. The actual prompts are provided in Figures 9 and 10 in Appendix B. We then define the milestone achievement count for a trace $\gamma$ as:
@@ 5.1 Experiment Setup
We consider two distinct agent frameworks in this work. Following whoandwhen, we begin by conducting experiments using Magentic-One (M1) Magentic-One, a popular LLM-based multi-agent framework that was also used for collecting failure traces in WW. We enable DoVer on M1 by adapting the manual debugging tool AGDebugger AGDebuger so that DoVer can (without human intervention): (i) save the state at each step as a checkpoint; (ii) load checkpoints from a prior failed run; (iii) intervene by modifying an agent message at a specified step; and (iv) resume execution from the intervened step.
To evaluate DoVer’s generality, we further construct a MathChat multi-agent system using a second framework, AutoGen2 (AG2) autogen02; wu2023mathchat. This MathChat system is instantiated using the prompts provided in MAST MAST, and we extend AG2 with checkpointing and re-execution capabilities analogous to those in M1. Implementation details and practical lessons for reducing the integration burden on new frameworks are provided in Appendix C.
@@
We first follow whoandwhen and include all cases in the Hand Crafted category of WW: 28 cases from AssistantBench AssistantBench and 29 cases spanning all three GAIA levels GAIA. To increase data volume, we additionally include all 53 Level-1 cases from GAIA’s validation set; after excluding those already present in WW, this yields 45 extra cases. We refer to these three sets as WW-AB, WW-GAIA, and GAIA-Level-1, respectively.
For the MathChat system, following MAST MAST, we additionally use the GSMPlus dataset li2024gsm. Concretely, we adopt the 2,400 examples in the “testmini” split and re-collect execution traces with checkpoints for all problems, forming the GSMPlus setting used in our AG2-based experiments.
@@
We begin with an initial run over all cases and evaluate outcomes to identify failure traces. The failure logs published with WW and MAST are not directly usable for our purposes: logs of agent messages alone are insufficient to support replay and targeted intervention. The number of failed cases per dataset is given in Table 1 (“Failed Cases” column). The overall success rate on the full GAIA Level-1 set matches the value reported for Magentic-One Magentic-One, suggesting that our collected failure traces faithfully reflect M1’s capabilities. As in WW and MAST, we use GPT-4o both to generate traces and to power the intervened session runs.
@@ 5.3 Ablation Study
To contextualize DoVer’s gains, we compare against two self-improvement style methods adapted from Self-Refine  madaan2023self and CRITIC  gou2023critic in the WW-GAIA setting. In the Self-Refine-style baseline, the underlying LLM first critiques the final answer given the full session log and then generates a revised answer in a second pass. In the CRITIC-style baseline, we similarly elicit feedback from the final answer and log, but inject this feedback as an additional agent message and allow the agent system to run for one extra round so that all tools and sub-agents can react to it. Across all 26 failed WW-GAIA cases, neither baseline is able to flip any failure into success (0% recovery), whereas DoVer recovers 17.6% of trials (Table  2). Further examination reveals why these self-improvement methods fall short: trial trajectories between the initial failure point and the final answer are often long, noisy, and highly divergent, making end-of-trace refinement insufficient for reliably redirecting the system. In contrast, DoVer performs in-situ interventions at potential failure points, enabling timely and targeted corrections that are essential in multi-agent settings.

### v3

@@ 1 Introduction
The advancement of Large Language Models (LLMs) has led to the rapid rise of LLM-based agent systems, particularly multi-agent architectures where agents of different roles work collaboratively to solve complex tasks Li et al. (2023); Wu et al. (2023a); Hong et al. (2024). As these systems are increasingly developed and deployed in production, the need to debug their failures becomes inevitable during their lifecycle. Importantly, by “failures” we do not refer to conventional software errors (e.g., exceptions or crashes), but rather to the errors where the system executes without interruption yet produces incorrect or unsatisfactory results Mialon et al. (2024); Yoran et al. (2024). Such failures frequently arise in scenarios where one diagnoses why an agent system underperforms on benchmark tasks during the development phase, or when one addresses user-reported dissatisfaction (e.g., a ‘thumbs-down’ signal with textual feedback) from an online deployed system.
Debugging failures in LLM-based agent systems presents unique challenges. These systems typically involve multiple rounds of LLM calls, each with extensive textual context, making manual log inspection labor-intensive. Furthermore, in multi-agent tasks, tracking inter-agent information flow is crucial, as failures often stem from Inter-Agent Misalignment Cemri et al. (2025). Recent efforts address these issues by using LLMs to analyze system failures Zhuge et al. (2024); Zhang et al. (2025c; a), often via single-agent, single-step failure attribution. In this method, an execution log is input to an LLM tasked with identifying the agent and step responsible for the failure Zhang et al. (2025c). However, as shown in our reproduction study (Section 3), log-based failure attribution is fundamentally limited by the uncertainty of ground-truth annotations. This uncertainty arises for several reasons: agent systems often employ multiple strategies (e.g., ReAct Yao et al. (2023)) with distinct failure points in a single session, and inter-agent misalignment can render the assignment of responsibility to a single agent or step ambiguous.
@@
DoVer also supports interventions across multiple steps rather than restricting to single-point edits. By decomposing the failure trace into separate trials, we intervene at each and assess the impact. Within the Magnetic-One (M1) Fourney et al. (2024) agent framework, our experiments show that DoVer recovers 18% and 28% of failures on datasets from AssistantBench Yoran et al. (2024) and GAIA Mialon et al. (2024), respectively. It further enables the validation or refutation of 30–60% of failure hypotheses, depending on task complexity. To demonstrate the generality of DoVer, we apply it to another agent framework, AG2 AutoGen2 (2025), using a different dataset, GSMPlus Li et al. (2024). DoVer again performs effectively, achieving a 49% flip rate on failure cases.
@@ 2 Related Work
LLM-based agent systems exhibit diverse and frequent failure patterns that accumulate along long execution logs. To characterize why and where errors arise, MAST Cemri et al. (2025) catalogs failures across task interpretation, planning, tool/environment interaction, and verification, while evaluation frameworks Zhuge et al. (2024); Arabzadeh et al. (2024) argue that end-to-end pass/fail is too coarse and introduce requirement graphs or task-utility criteria that reveal where progress stalls. Extending this perspective to execution logs, TRAIL Deshpande et al. (2025) creates turn-level traces and a fine-grained taxonomy (reasoning, planning, execution), empirically showing that even strong long-context models struggle at trace debugging.
A parallel line of work seeks failure attribution: identifying the earliest decisive step or agent that is responsible for the earliest sufficient cause of failure Zhang et al. (2025c; a). However, these attributions are inferred from logs and remain an untested hypothesis unless validated by execution. In DoVer, we treat attribution as a hypothesis to be tested. We apply a targeted edit at the implicated location (message, plan, tool call), and rerun the system, judging success by milestone/utility gains. This places emphasis on verified repair and is consistent with trajectory-aware evaluation advocated in Arabzadeh et al. (2024); Deshpande et al. (2025).
Recent contemporaneous studies [^1] These works appeared during the final stage of this work preparation or the subsequent peer-review period. further advance failure attribution by introducing reasoning-driven judges Zhu et al. (2025a), abduct–act–predict counterfactual scaffolding West et al. (2025), causal-inference formulations Ma et al. (2025), spectrum-based failure attribution approach Ge et al. (2025), hierarchical error-attribution method Banerjee et al. (2025), and graph-guided failure tracing approach Zhang et al. (2025b). Notably, the observation in West et al. (2025) that adding explicit step indices improve attribution aligns with our own prompt-refinement study in Section 3.
Emerging works have also explored specialized failure-tracer models trained on curated success/failure trajectories Zhang et al. (2025a); Kong et al. (2025). These approaches are orthogonal to DoVer and could be incorporated into the initial failure-attribution stage of DoVer to strengthen its diagnostic signal.
@@ 2.2 Debugging Approaches for LLM-based Agent Systems
Beyond failure analysis and attribution, several systems explore how to debug trajectories, i.e., intervention and replay. Human-in-the-loop tools such as AGDebugger Epperson et al. (2025) enable rewind/edit/re-execute with trace visualization, and graph runtimes like LangGraph LangChain (2025) provide checkpoints, interrupts, and “time-travel” branching. These demonstrate that small, targeted interventions often work, but they are manual and hard to scale. The recent AgentDebug Zhu et al. (2025b) work employs an intervention-driven methodology similar to DoVer, although it does not place particular emphasis on multi-agent settings.
From the software-repair perspective, Rahardja et al. (2025) package real agent-system issues into executable environments with failing tests (AgentIssue-Bench) and find low resolution rates for current software engineering agents, highlighting the difficulty of maintaining agent software. Industrial experience at Google similarly evaluates agent-based program repair on production bugs, showing promise but also current limits Rondon et al. (2025). This motivates auto debugging that verifies edits within the original run and measures intermediate progress. DoVer does so by choosing a minimal intervention, re-running the trajectory, and scoring milestone/utility gains, in line with Zhuge et al. (2024); Deshpande et al. (2025).
@@ 3 From Log-Based Attribution to Intervention-Based Debugging
A common setup for log-based failure attribution proceeds as follows. The input is a session log, typically from an LLM-driven multi-agent system, covering the full trajectory from an initial user query to either a final answer or termination caused by system-defined stopping rules (e.g., a maximum number of replanning rounds). The log is a sequence of agent messages organized by turns. [^2] We do not consider logs from asynchronous agent executions in this work. The task is to identify the earliest single agent and single step responsible for the failure. A failure step is defined as a decisive error: if one were to replace the agent’s incorrect action at that step with a correct one, the trajectory would subsequently succeed. To evaluate this setup, the Who&When (WW) dataset was introduced in Zhang et al. (2025c). In WW, multiple annotators independently reviewed session logs and then reconciled discrepancies through discussion. Reported results show that state-of-the-art LLMs at the time achieved below 10% accuracy on failure step attribution.
@@
Figure 1: Failure trace of Case 3 in WW-HC, illustrating ambiguity in failure attribution. The session consists of four distinct trials, each initiated by a plan update and executed via a ReAct-style Yao et al. (2023) loop. Different strategies (e.g., direct scrolling in Trial 1 vs. calendar navigation in Trial 2) yield separate error points, making single-step attribution across the session inherently ambiguous. Trial 2 (Steps 53–55) further shows inter-agent misalignment: the Orchestrator issued an invalid instruction, while the WebSurfer compounded the error by executing an unrelated action.
@@
(1) Multiple trials within a single session. Modern agentic systems frequently employ ReAct-style Yao et al. (2023) loops with repeated planning–execution cycles, producing multiple trials of task solving. We define a trial as the contiguous span starting from a planning step and continuing through the execution steps for that plan. Each trial may contain its own decisive error step(s), especially when exploring new strategies or branches. For instance, Figure 1 illustrates the failure trace of Case 3 in WW-HC, which comprises four distinct trials within a single session. Different strategies were explored (e.g., Trial 1 attempted to locate the target webpage via direct scrolling, whereas Trial 2 used the calendar feature), each introducing distinct potential error points. Consequently, enforcing a single-step attribution across the entire session is intrinsically ambiguous. In our review, 9 of the 14 uncertain cases exhibit this pattern.
(2) Ambiguity from multi-agent coordination. Attribution to a specific agent/step can be unclear when the underlying issue stems from inter-agent coordination. For instance, when a sub-agent fails to carry out an instruction, responsibility may lie with the sub-agent’s capabilities or with ambiguous/misaligned guidance from a supervising agent. In Trial 2 of Figure 1, for example, the Orchestrator agent instructed the WebSurfer agent to click on a non-existent control. Rather than reporting the issue, the WebSurfer instead clicked on an unrelated control that had no connection to the target year “2015”. In this case, both agents exhibited failures, making attribution to either one alone inappropriate. Such ambiguous attribution aligns with the Inter-Agent Misalignment category identified in Cemri et al. (2025). We observed this phenomenon in 5 of the 14 uncertain cases.
@@ 4.1 DoVer Pipeline
Trial Segmentation. After a task is executed by an agentic system, we have obtained a long execution session log $\tau=\{(a_{t},m_{t},\sigma_{t})\}_{t=1}^{T}$, where $a_{t}$ represents the active agent that produces the message $m_{t}$ at step $t$, and $\sigma_{t}$ keeps stateful information that is necessary for state restore and replay. For example, this includes historical context (e.g., prior messages sent to LLMs) as well as browsing history for agents acting in a web-browsing role. Since modern LLM-based agent systems Fourney et al. (2024); SmolAgents (2025); AutoGen2 (2025) have the self-reflection capabilities Yao et al. (2023); Shinn et al. (2023), and re-plan after reflection, we then segment the trace $\tau$ into trials $\tau^{i}$} using re-plan steps as segmentation point, as illustrated in Figure 2. This trial segmentation shortens context so LLMs can reason about a single causal chain, and enables independent and parallel interventions for efficiency. Thus, multiple hypotheses can be proposed for a single session trace, which is essential as many failures admit more than one viable repair.
@@ 4.2 Evaluation Metrics
Specifically, when an intervention does not yield full success and human-annotated solution steps are available (e.g., in the GAIA benchmark Mialon et al. (2024)), we could evaluate the degree of progress by comparing the new execution trace against human-annotated steps. For each log $\tau$, we extract up to $K\leq 5$ milestones $\{\mathbf{m}^{k}\}_{k=1}^{K}$ and measure how many of these milestones the new trace accomplishes. Both the extraction of milestones and their evaluation are performed using LLMs with carefully designed instructions. The actual prompts are provided in Figures 9 and 10 in Appendix B. We then define the milestone achievement count for a trace $\gamma$ as:
@@ 5.1 Experiment Setup
We consider two distinct agent frameworks in this work. Following Zhang et al. (2025c), we begin by conducting experiments using Magentic-One (M1) Fourney et al. (2024), a popular LLM-based multi-agent framework that was also used for collecting failure traces in WW. We enable DoVer on M1 by adapting the manual debugging tool AGDebugger Epperson et al. (2025) so that DoVer can (without human intervention): (i) save the state at each step as a checkpoint; (ii) load checkpoints from a prior failed run; (iii) intervene by modifying an agent message at a specified step; and (iv) resume execution from the intervened step.
To evaluate DoVer’s generality, we further construct a MathChat multi-agent system using a second framework, AutoGen2 (AG2) AutoGen2 (2025); Wu et al. (2023b). This MathChat system is instantiated using the prompts provided in MAST Cemri et al. (2025), and we extend AG2 with checkpointing and re-execution capabilities analogous to those in M1. Implementation details and practical lessons for reducing the integration burden on new frameworks are provided in Appendix C.
@@
We first follow Zhang et al. (2025c) and include all cases in the Hand Crafted category of WW: 28 cases from AssistantBench Yoran et al. (2024) and 29 cases spanning all three GAIA levels Mialon et al. (2024). To increase data volume, we additionally include all 53 Level-1 cases from GAIA’s validation set; after excluding those already present in WW, this yields 45 extra cases. We refer to these three sets as WW-AB, WW-GAIA, and GAIA-Level-1, respectively.
For the MathChat system, following MAST Cemri et al. (2025), we additionally use the GSMPlus dataset Li et al. (2024). Concretely, we adopt the 2,400 examples in the “testmini” split and re-collect execution traces with checkpoints for all problems, forming the GSMPlus setting used in our AG2-based experiments.
@@
We begin with an initial run over all cases and evaluate outcomes to identify failure traces. The failure logs published with WW and MAST are not directly usable for our purposes: logs of agent messages alone are insufficient to support replay and targeted intervention. The number of failed cases per dataset is given in Table 1 (“Failed Cases” column). The overall success rate on the full GAIA Level-1 set matches the value reported for Magentic-One Fourney et al. (2024), suggesting that our collected failure traces faithfully reflect M1’s capabilities. As in WW and MAST, we use GPT-4o both to generate traces and to power the intervened session runs.
@@ 5.3 Ablation Study
To contextualize DoVer’s gains, we compare against two self-improvement style methods adapted from Self-Refine  Madaan et al. (2023) and CRITIC  Gou et al. (2023) in the WW-GAIA setting. In the Self-Refine-style baseline, the underlying LLM first critiques the final answer given the full session log and then generates a revised answer in a second pass. In the CRITIC-style baseline, we similarly elicit feedback from the final answer and log, but inject this feedback as an additional agent message and allow the agent system to run for one extra round so that all tools and sub-agents can react to it. Across all 26 failed WW-GAIA cases, neither baseline is able to flip any failure into success (0% recovery), whereas DoVer recovers 17.6% of trials (Table  2). Further examination reveals why these self-improvement methods fall short: trial trajectories between the initial failure point and the final answer are often long, noisy, and highly divergent, making end-of-trace refinement insufficient for reliably redirecting the system. In contrast, DoVer performs in-situ interventions at potential failure points, enabling timely and targeted corrections that are essential in multi-agent settings.